# Smoothing of Headland Path Edges and Headland-to-Mainfield Lane Transitions Based on a Spatial Domain Transformation and Linear Programming

### v1

@@ 4.1 Extending Discussion
Following similarly to Oh et al. (2023) an analytic approach and as earlier work within the agricultural context, Sabelhaus et al. (2013) is briefly discussed. They present analytic formulas and visualisations for 9 different symmetric lane-to-lane transitions (Omega-turn, U-turn, Gap-turn, Slope-turn,, Fishtail-turn,, Fishtail-turn with slope, Minimal longitudinal width, Pinhole-turn, Reversal pinhole turn) for continuous-curvature path planning. They also present three simulation experiments with longitudinal, lateral and angle offset of start and end point. Characteristically, all paths are constructed based on path segments with maximum curvature rate (clothoids with so called maximal sharpness) and circle segments with maximum curvature. Five comments are made. First, the methods are not suitable for Problem 1 since for that problem paths are needed that are not always and necessarily operating at the maximum curvature rate limit. Second and interestingly, in general none of the 9 cases can handle Problem 2 (e.g. for transitions as shown in Fig. 5). The Minimal longitudinal width-case resembles a transition as shown in Fig. 5, but it has a specific condition $d=R_{\text{big}}$ (Sect. 2.3.7 in Sabelhaus et al. (2013)) that limits its usage. (Another reason herefore is that Sabelhaus et al. (2013) focus only on lane-to-lane transitions and not dicuss headland-to-lane transitions.) Third, Sabelhaus et al. (2013) acknowledge that ’manoeuvres with start or end curvature $\kappa\neq 0$ are not feasible with continuous curvature turns’. In contrast, for proposed LP-based method this case can be handled, by adding an initial input constraint $\delta_{0}=\delta_{\text{start}}$ or $\delta_{N-1}=\delta_{\text{end}}$ to (6) or (7) for Problem 1 or 2, respectively, and by observing that curvature can be related to steering via $\kappa=\frac{\text{tan}(\delta)}{l}$ (for the assumption of no-slip conditions of the wheels and instantaneous curvature of the rear axle). This is made possible since the LP-based approach permits to incorporate state and input constraints in a structured manner. Fourth, as discussed for Oh et al. (2023), likewise the approach of Sabelhaus et al. (2013) is limited to a specific vehicle model (no slip condition). Fifth, the proposed spatial LP-method can also handle lane-to-lane transitions. For illustration see Fig. 10 with multiple non-symmetric start end points and an interlane-distance of 5m. The simplicity of proposed method is underlined. For lane-to-lane transitions a Dubins-path is computed between the intersection points of headland path and mainfield lanes. The Dubins path is then used as reference path in the second hierarchical step solving a spatial-based LP to obtain a refined path. This paragraph underlines the flexibility of proposed method, in particular for shaping solutions to desired behavior.
@@
A brief comment about path planning methods by neural networks is made (see a simplistic algorithm in Graf Plessen (2020) and the references therein). The key argument against the employment of neural networks is that both path planning Problems 1 and 2 deal with an infinite number of possible tasks for headland path edges and headland-to-mainfield lane transitions and require a continuous action space. In order to encode a path planning module in a neural network and enormous number of training examples and their ground-truth solutions would be required, since there is an enormous variety in nature of different field contours, which would make a large training set mandatory. To underline the continous control aspect, exemplarily see Plappert et al. (2018) where policies for continuous control are learned. However, each of the evaluation problems receives a reward either (i) upon moving forward (HalfCheetah, Hopper, Swimmer, Walker2D, SparseHalfCheetah, SwimmerGather) or (ii) upon reaching a quasi-stationary goal position (InvertedDoublePendulum, InvertedPendulum, Reacher, SparseCartpoleSwingup, SparseMountainCar, SparseDoublePendulum). There is no option for shaping of paths by accounting for precision constraints. Training time and hardware requirements are not stated. In contrast, proposed method is based on linear programming, a hyperparameter-free objective function, very fast solve times, and can incorporate multivariate nonlinear vehicle dynamics and constraints on both states as well as control inputs in a structured manner to allow flexibility and generate deterministic precision paths, which is required for both Problems 1 and 2.

### v3

@@ 5 Discussion
Following similarly to Oh et al. (2023) an analytic approach and as earlier work within the agricultural context, Sabelhaus et al. (2013) is briefly discussed. They present analytic formulas and visualisations for 9 different symmetric lane-to-lane transitions (Omega-turn, U-turn, Gap-turn, Slope-turn, Fishtail-turn, Fishtail-turn with slope, Minimal longitudinal width, Pinhole-turn, Reversal pinhole turn) for continuous-curvature path planning. They also present three simulation experiments with longitudinal, lateral and angle offset of start and end point. Characteristically, all paths are constructed based on path segments with maximum curvature rate (clothoids with so called maximal sharpness) and circle segments with maximum curvature. Five comments are made. First, the methods are not suitable for Problem 1 since for that problem paths are needed that are not always and necessarily operating at the maximum curvature rate limit. Second and interestingly, in general none of the 9 cases can handle Problem 2 (e.g. for transitions as shown in Fig. 6). The Minimal longitudinal width-case resembles a transition as shown in Fig. 6, but it has a specific condition $d=R_{\text{big}}$ (Sect. 2.3.7 in Sabelhaus et al. (2013)) that limits its usage. (Another reason herefore is that Sabelhaus et al. (2013) focus only on lane-to-lane transitions and not dicuss headland-to-lane transitions.) Third, Sabelhaus et al. (2013) acknowledge that ’manoeuvres with start or end curvature $\kappa\neq 0$ are not feasible with continuous curvature turns’. In contrast, for proposed LP-based method this case can be handled, by adding an initial input constraint $\delta_{0}=\delta_{\text{start}}$ or $\delta_{N-1}=\delta_{\text{end}}$ to (6) or (7) for Problem 1 or 2, respectively, and by observing that curvature can be related to steering via $\kappa=\frac{\text{tan}(\delta)}{l}$ (for the assumption of no-slip conditions of the wheels and instantaneous curvature of the rear axle). This is made possible since the LP-based approach permits to incorporate state and input constraints in a structured manner. Fourth, as discussed for Oh et al. (2023), likewise the approach of Sabelhaus et al. (2013) is limited to a specific vehicle model (no slip condition). Fifth, the proposed spatial LP-method can also handle lane-to-lane transitions. For illustration see Fig. 11 with multiple non-symmetric start and end points and an interlane-distance of 5m. The simplicity of proposed method is underlined. For lane-to-lane transitions a Dubins-path is computed between the intersection points of headland path and mainfield lanes. The Dubins path is then used as reference path in the second hierarchical step solving a spatial-based LP to obtain a refined path. This paragraph underlines the flexibility of proposed method, in particular for shaping solutions to desired behavior.
@@
A brief comment about path planning methods by neural networks is made (see a simplistic algorithm in Graf Plessen (2020) and the references therein). The key argument against the employment of neural networks is that both path planning Problems 1 and 2 deal with an infinite number of possible tasks for headland path edges and headland-to-mainfield lane transitions and require a continuous action space. To encode a path planning module in a neural network an enormous number of training examples and their ground-truth solutions would be required, since there is an enormous variety in nature of different field contours, which would make a large training set mandatory. To underline the continous control aspect, exemplarily see Plappert et al. (2018) where policies for continuous control are learned. However, each of the evaluation problems receives a reward either (i) upon moving forward (HalfCheetah, Hopper, Swimmer, Walker2D, SparseHalfCheetah, SwimmerGather), or (ii) upon reaching a quasi-stationary goal position ( InvertedDoublePendulum, InvertedPendulum, Reacher, SparseCartpoleSwingup, SparseMountainCar, SparseDoublePendulum). There is no option for path shaping by accounting for precision constraints. Training time and hardware requirements are not stated. In contrast, proposed method is based on linear programming, with hyperparameter-free objective function, fast solve times, and can incorporate multivariate nonlinear vehicle dynamics and constraints on both states and control inputs in a structured manner to allow flexibility and generate deterministic precision paths, which is required for both Problems 1 and 2.